### v1

# Density Frankl–Rödl on the Sphere

Abstract
We establish a density variant of the Frankl–Rödl theorem on the sphere $\mathbb{S}^{n-1}$, which concerns avoiding pairs of vectors with a specific distance, or equivalently, a prescribed inner product. In particular, we establish lower bounds on the probability that a randomly chosen pair of such vectors lies entirely within a measurable subset $A\subseteq\mathbb{S}^{n-1}$ of sufficiently large measure. Additionally, we prove a density version of spherical avoidance problems, which generalize from pairwise avoidance to broader configurations with prescribed pairwise inner products. Our framework encompasses a class of configurations we call inductive configurations, which include simplices with any prescribed inner product $-1<r<1$. As a consequence of our density statement, we show that all inductive configurations are sphere Ramsey.
Keywords: Frankl–Rödl, Sphere Ramsey, Sphere Avoidance, Reverse Hypercontractivity, Forbidden Angles
1 Introduction
The Frankl–Rödl theorem [FR87] is a foundational result in extremal combinatorics and theoretical computer science. It states that for any fixed $10<\gamma<$ 1, assuming $(1-\gamma)n$ is even, if a set $A\subseteq\{-1,1\}^{n}$ contains no pair of distinct points at Hamming distance exactly $(1-\gamma)n$, then the fractional size of $A$ must be exponentially small. That is, there exists a constant $\epsilon=\epsilon(\gamma)<1$ such that
$|A|/2^{n}\leq\epsilon^{n}.$
The corresponding Frankl–Rödl graph ${\mathrm{FR}_{\gamma}^{n}}$ is defined on vertex set $\{-1,1\}^{n}$ with edges between points at Hamming distance $(1-\gamma)n$, with the Frankl-Rödl theorem bounding the independence number of this graph. This theorem has found broad applications in theoretical computer science, particularly in the analysis of hardness of approximation. The graph ${\mathrm{FR}_{\gamma}^{n}}$ has been used to construct integrality gap instances for problems such as 3-Coloring [KG98, KMS98, Cha02, AG11, KOTZ14] and Vertex-Cover [KG98, Cha02, ABL02, ABLT06, GMT08, GMPT10, KOTZ14].
The Frankl–Rödl theorem extends naturally from pairs of points to general forbidden configurations. A configuration of $k$ vertices is specified by a set of pairwise distances (or equivalently, inner products). Frankl and Rödl [FR87] showed that for any fixed $k$, any subset $A\subseteq\{-1,1\}^{4n}$ that avoids $r$ pairwise orthogonal vectors must also have exponentially small fractional size. This result has been applied to show an $\Omega(\log n)$ integrality gap for the SDP relaxation of Min-Multicut [ACMM05].
While the original theorem asserts that for any $A\subseteq\{-1,1\}^{n}$ of size $|A|/2^{n}\geq\epsilon(\gamma)^{n}$,
$\mathop{\bf Pr\/}_{\begin{subarray}{c}x,y\in\{-1,1\}^{n}\\ x\cdot y=(2\gamma-1)n\end{subarray}}(x\in A,y\in A)>0,$
it is natural to ask whether one can give a quantitative lower bound on this probability in terms of the density $|A|/2^{n}$. Benabbas, Hatami, and Magen [BHM12] answered this by proving a density version of the Frankl–Rödl theorem: for any $10<\alpha<$ 1, if $|A|/2^{n}\geq\alpha$, then
$\mathop{\bf Pr\/}_{\begin{subarray}{c}x,y\in\{-1,1\}^{n}\\ x\cdot y=(2\gamma-1)n\end{subarray}}(x\in A,y\in A)\geq 2\left(\alpha/2\right)^{\frac{2}{1-|2\gamma-1|}}-o(1).$
This result is based on the reverse hypercontractivity of the Bonami–Beckner semigroup [MOR+06], and was used to prove integrality gaps for Vertex-Cover. Later Kauers et. al [KOTZ14] showed that the proof can be conducted in the sum-of-squares (SOS) proof system. In particular, they showed that for any $40<\gamma\leq 1/$ 4, the SOS/Lasserre SDP hierarchy at degree $4\left\lceil\frac{1}{4\gamma}\right\rceil$ certifies that the maximum independent set in ${\mathrm{FR}_{\gamma}^{n}}$ has fractional size $o(1)$. This implies that a degree-4 algorithm from the SOS hierarchy can certify that the ${\mathrm{FR}_{\gamma}^{n}}$ SDP integrality gap instances for 3-Coloring have chromatic number $\sigma(1)$, and that a degree-$\left\lceil 1/\gamma\right\rceil$ SOS algorithm can certify the ${\mathrm{FR}_{\gamma}^{n}}$ SDP integrality gap instances for Min-Vertex-Cover has minimum vertex cover $>1-o(1)$.
In the continuous setting, the sphere $\mathbb{S}^{n-1}$ provides a natural high-dimensional geometric analogue. This motivates the study of spherical avoidance problems: how dense can a subset $A\subseteq\mathbb{S}^{n-1}$ be while avoiding a fixed configuration of pairwise inner products? Let $\Delta_{k}(n,r)$ denote the $k$-simplex in $\mathbb{S}^{n-1}$ with pairwise inner product $r$. Witsenhausen’s problem [Wit74] asks for the maximum density of a measurable set $A\subseteq\mathbb{S}^{n-1}$ avoiding orthogonal pairs, i.e., $\Delta_{2}(n,0)$. Frankl and Wilson [FW81] gave the first exponentially decreasing upper bound $\sigma(A)\leq(1+o(1))(1.13)^{-n}$, where $\sigma$ denotes the uniform surface measure on $\mathbb{S}^{n-1}$. Kalai [KW09] conjectured that the extremal set consists of two opposite caps of geodesic radius $\pi/4$; this Double Cap Conjecture implies new lower bounds for the measurable chromatic number of $\mathbb{R}^{n}$ [DP16].
Regev and Klartag [RK11] established a density version of the Frankl–Rödl theorem on the sphere for pairs of orthogonal vectors:
$\mathop{\bf Pr\/}_{\begin{subarray}{c}x,y\in{\mathbb{S}^{n-1}}\\ x\cdot y=0\end{subarray}}(x\in A,\ y\in A)\geq 0.9\,\sigma(A)^{2},$
valid for any measurable set $A\subseteq{\mathbb{S}^{n-1}}$ with $\sigma(A)\geq C\exp(-cn^{1/3})$, where $C$ and $c$ are universal constants. This result was a key component in their proof of an $\Omega(n^{1/3})$ lower bound on the classical communication complexity of the Vector in Subspace Problem (VSP). As a major consequence, they resolved a long-standing open question posed by Raz [Raz99], demonstrating that quantum one-way communication is indeed exponentially stronger than classical two-way communication.
For configurations $\Delta_{k}(n,r)$ with $r>0$, Castro-Silva et al. [CSdOFSV22] proved that for any $k\geq 2$, there exists $c=c(k,r)<1$ such that any $A\subseteq\mathbb{S}^{n-1}$ avoiding $\Delta_{k}(n,r)$ satisfies $\sigma(A)\leq(c+o(1))^{n}$. A configuration is sphere Ramsey if any $c$-coloring of of the sphere, there exists a monochromatic congruent copy of the configuration. Matoušek and Rödl [MR95] showed that any configuration $P$ with circumradius less than 1 is sphere Ramsey.
This sphere Ramsey theorem was recently used by Brakensiek, Guruswami, and Sandeep [BGS23] to demonstrate integrality gaps for the basic SDP relaxation of certain promise CSPs: namely Boolean symmetric promise CSPs defined by a single predicate pair that lack Majority or Alternate-Threshold (AT) polymorphisms of all odd arities. Via Raghavendra’s general connection tightly linking SDP integrality gaps to Unique-Games hardness [Rag08], this enabled BGS to conclude that such promise CSPs do not admit a robust satisfiability algorithm (in the sense of Zwick [Zwi98]) under the Unique Games conjecture. Complemeting this hardness result, BGS gave a robust satisfiability algorithm for Boolean promise CSPs that admit Majority polymorphisms of all odd arities or AT polymorphisms of all odd arities—the algorithm applied in the general promise CSP setting that could have multiple predicate pairs. Together these results led to a dichotomy theorem with respect to robust satisfiability for Boolean symmetric promise CSPs defined by a single predicate. Towards extending their hardness result to Boolean symmetric PCSPs that could include multiple predicate pairs, BGS posed the following problem: can one show a density version of spherical avoidance problems for the configuration $(x_{1},\cdots,x_{b},-x_{b+1},\cdots,-x_{k})$ where $(x_{1},\cdots,x_{b},x_{b+1},\cdots,x_{k})\in\Delta_{k}(n,r)$. Namely, obtain a non-trivial lower bound on the quantity
$\mathop{\bf Pr\/}_{(x_{1},\cdots,x_{b},x_{b+1},x_{k})\in\Delta_{k}(n,r)}\left(x_{1}\in A,\cdots,x_{b}\in A,-x_{b+1}\in A,\cdots,-x_{k}\in A\right).$
in terms of $\sigma(A)$. While we do not directly resolve the exact configuration posed by BGS, our result applies to a broad class of configurations that includes closely related structures.
In this paper, we show that for any $-\frac{1}{k-1}<r<1$, there exists a constant $C=C(r)$ such that
$\mathop{\bf Pr\/}_{x_{1},\dots,x_{k}\in\Delta_{k}(n,r)}(x_{i}\in A\ \forall i)\geq\Omega_{k,r}\left(\sigma(A)^{C}\right)$
for all measurable $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega(n^{-2/3})$. Our result generalizes the work of Castro-Silva et al. to simplices with any $-1<r<1$, and more broadly, to a class of inductive configurations defined in Section 4. As a consequence, we show that all inductive configurations are sphere Ramsey.
Inspired by the techniques of Benabbas et al. [BHM12], who used reverse hypercontractivity of the Bonami–Beckner semigroup to show the density variant of Frankl–Rödl on $\{-1,1\}^{n}$, we develop analogous tools on the sphere. Specifically, we prove reverse hypercontractivity of the operator
$A_{t}f(x):=\mathop{\mathbb{E}\/}_{\begin{subarray}{c}y\in\mathbb{S}^{n-1}\\ x\cdot y=e^{-t}\end{subarray}}[f(y)],$
which enables us to derive density versions of spherical avoidance for inductive configurations. In Section 2, we analyze the eigen-decomposition of $A_{t}$; in Section 3, we relate $A_{t}$ to the Poisson Markov semigroup $P_{t}$ and establish reverse hypercontractive inequalities. Finally, in Section 4, we prove our main results on density versions of the Frankl–Rödl theorem on the sphere and density spherical for inductive configurations.
2 Eigen-decomposition via Spherical Harmonics
In order to analyze the eigen-decomposition of $A_{t}$, we first introduce some fundamental notions from the theory of spherical harmonics. This decomposition of $A_{t}$ plays a central role in our analysis of reverse-hypercontractive. For a comprehensive background, we refer to the standard references [Mül06, SW71] and notes such as [Gal09].
2.1 Spherical Harmonics
For each integer $n\geq 2$, let $\mathbb{S}^{n-1}\subseteq\mathbb{R}^{n}$ denote the unit $n$-sphere, defined by
$\mathbb{S}^{n-1}=\{x\in\mathbb{R}^{n}:x\cdot x=1\}.$
Given $p>0$ and a measurable function $f:\mathbb{S}^{n-1}\to\mathbb{R}$, we define its $L^{p}$ norm by
[…]
We focus in particular on $L^{2}(\mathbb{S}^{n-1})$, the space of square-integrable functions. It admits an orthonormal basis consisting of spherical harmonics $\{S_{k,\ell}\}$, indexed by integers $k\geq 0$ (the degree) and $1\leq\ell\leq c_{k,n}$, where
$c_{k,n}=\binom{k+n-2}{k}+\binom{k+n-3}{k-1}.$
Each $S_{k,\ell}$ is the restriction to $\mathbb{S}^{n-1}$ of a harmonic, homogeneous polynomial of degree $k$ on $\mathbb{R}^{n}$. We denote by $\mathcal{S}_{k}\subseteq L^{2}(\mathbb{S}^{n-1})$ the finite-dimensional subspace spanned by spherical harmonics of degree $k$:
$\mathcal{S}_{k}:=\operatorname{span}\{S_{k,1},\dotsc,S_{k,c_{k,n}}\}.$
Notably, $\mathcal{S}_{0}$ consists of the constant functions. Rotations $U\in\mathrm{SO}(n)$ act on $f\in L^{2}(\mathbb{S}^{n-1})$ via
$(Uf)(x):=f(U^{-1}x).$
Each space $\mathcal{S}_{k}$ is invariant under the action of $\mathrm{SO}(n)$, and furnishes an irreducible representation, i.e. any subspace $V\subseteq{\mathcal{S}_{k}}$ that is invariant under rotations must either be $\{0\}$ or ${\mathcal{S}_{k}}$. Moreover, for $n\geq 3$, the spaces $\mathcal{S}_{k}$ for different $k$ have different dimensions therefore are inequivalent as representations. Throughout the remainder of this paper, we assume $n\geq 3$.
For $t\geq 0$, define an operator $A_{t}:L^{2}(\mathbb{S}^{n-1})\to L^{2}(\mathbb{S}^{n-1})$ by
[…]
where $\sigma_{x,r}$ is the uniform probability measure on the $(n-2)$-subsphere $\{y\in\mathbb{S}^{n-1}:x\cdot y=r\}$.
The operator $A_{t}$ is easily seen to commute with rotations. It is also self-adjoint:
$\left\langle{A_{t}f,g}\right\rangle=\mathop{\mathbb{E}\/}_{x\in{\mathbb{S}^{n-1}}}[g(x)A_{t}f(x)]=\mathop{\mathbb{E}\/}_{x\in{\mathbb{S}^{n-1}}}\left[\mathop{\mathbb{E}\/}_{y:x\cdot y=e^{-t}}[g(x)f(y)]\right]=\mathop{\mathbb{E}\/}_{y\in{\mathbb{S}^{n-1}}}\left[\mathop{\mathbb{E}\/}_{x:x\cdot y=e^{-t}}[g(x)f(y)]\right]=\left\langle{f,A_{t}g}\right\rangle.$
for all $f,g\in L^{2}(\mathbb{S}^{n-1})$.
Using standard techniques from representation theory, we establish that the spaces $\mathcal{S}_{k}$ are eigenspaces of $A_{t}$:
Lemma 2.1.
For any $k\geq 0$ and $Y_{k}\in\mathcal{S}_{k}$, we have
$A_{t}(Y_{k})=\mu_{k,t}Y_{k},$
where $\mu_{k,t}=G_{k}(e^{-t})$, with $G_{k}:[-1,1]\to\mathbb{R}$ denoting the degree-$k$ Gegenbauer polynomial (see,e.g., [Mül06]), given by
$G_{k}(r)=\mathbb{E}\left[\left(r+iX_{1}\sqrt{1-r^{2}}\right)^{k}\right],$
where $X=(X_{1},\dotsc,X_{n-1})$ is uniformly distributed on $\mathbb{S}^{n-2}$ and $X_{1}$ denotes any fixed coordinate.
Proof.
By Schur’s Lemma, any linear operator commuting with the group action must act as a scalar on each irreducible representation. Since $A_{t}$ commutes with rotations and $\mathcal{S}_{k}$ are inequivalent irreducible representations for different $k$, the restriction of $A_{t}$ to $\mathcal{S}_{k}$ must be a scalar multiple of the identity.
To identify the eigenvalue, consider the function $f_{k}(x):=G_{k}(x\cdot v)$ for some fixed $v\in\mathbb{S}^{n-1}$. It is known that $f_{k}\in\mathcal{S}_{k}$ [Mül06]. Then
$(A_{t}f_{k})(v)=G_{k}(e^{-t}),\quad\text{and}\quad f_{k}(v)=G_{k}(1)=1,$
showing that $\mu_{k,t}=G_{k}(e^{-t})$. ∎
2.2 Eigenvalue Estimates
We now estimate the eigenvalues $\mu_{k,t}$ for different regimes of $k$:
Lemma 2.2.
For all $0\leq k\leq n^{1/3}$, as $n\to\infty$, the eigenvalues satisfy
$|\mu_{k,t}-e^{-kt}|=O_{t}(n^{-1/3}),$
where the implicit constant depends only on $t$ but is independent of $k$.
Proof.
Let us first analyze the $k$-th moments of $X_{1}$. Notice that the density of $X_{1}$ is supported in $[-1,1]$ and proportional to $(1-x^{2})^{(n-4)/2}$. Since the density $X_{1}$ is an odd function, the odd moments vanishes, $\mathop{\mathbb{E}\/}[X_{1}^{k}]=0$ for odd $k$. For even $k$, there is an estimate
$\mathop{\mathbb{E}\/}[X_{1}^{k}]\leq\left(\frac{k}{n-4}\right)^{k/2}$
given by [RK11] Lemma 5.5. A direct computation gives
$\mu_{k,t}$ $=G_{k}(e^{-t})$
$=\ \mathop{\mathbb{E}\/}\left[\left(e^{-t}+iX_{1}\sqrt{1-e^{-2t}}\right)^{k}\right]$
$=\ \sum_{a=0}^{k}\left(i\sqrt{1-e^{-2t}}\right)^{a}e^{-t(k-a)}E[X_{1}^{a}]$
$=\ \sum_{a=0}^{\left\lfloor k/2\right\rfloor}(-1)^{a}(1-e^{-2t})^{a}e^{-t(k-2a)}E[X_{1}^{2a}]$
$=\ e^{-kt}+\sum_{a=1}^{\left\lfloor k/2\right\rfloor}(-1)^{a}(1-e^{-2t})^{a}e^{-t(k-2a)}E[X_{1}^{2a}]\ .$
Therefore
$|\mu_{k,t}-e^{-kt}|\leq{}$ $\sum_{a=1}^{\left\lfloor k/2\right\rfloor}(1-e^{-2t})^{a}e^{-t(k-2a)}\left(\frac{2a}{n-4}\right)^{a}$
[…]
Lemma 2.3.
For all $k\geq n^{1/3}$, as $n\to\infty$, the eigenvalues satisfy
$|\mu_{k,t}|=O\left(\frac{1}{n}\right).$
Proof.
Since $\mathop{\mathbb{E}\/}[X_{1}^{2}]\leq\frac{2}{n-4}$, by Markov’s inequality,
$\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})=\mathop{\bf Pr\/}(|X_{1}|^{2}\geq\frac{1}{4})\leq\frac{8}{n-4},$
which leads us to
$|\mu_{k,t}|$ $=\ |G_{k}(e^{-t})|$
$=\ \left|\mathop{\mathbb{E}\/}\left[\left(e^{-t}+iX_{1}\sqrt{1-e^{-2t}}\right)^{k}\right]\right|$
$\leq\ \mathop{\mathbb{E}\/}\left[\left|e^{-t}+iX_{1}\sqrt{1-e^{-2t}}\right|^{k}\right]$
$\leq\ \mathop{\mathbb{E}\/}\left[\left(e^{-2t}+(1-e^{-2t})|X_{1}|\right)^{k}\right]$
$\leq\ \mathop{\bf Pr\/}(|X_{1}|<\frac{1}{2})\left(e^{-2t}+\frac{1}{2}(1-e^{-2t})\right)^{k}+\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})$
$=\ O\left(\left(\frac{1+e^{-2t}}{2}\right)^{n^{1/3}}\right)+O(\frac{1}{n})\ =\ O(\frac{1}{n})$
as $n\rightarrow\infty$. ∎
3 Reverse Hypercontractivity of $A_{t}$
Hypercontractive inequalities play a central role in modern analysis, probability, and theoretical computer science, quantifying how semigroups “smooth out” irregularities over time by contracting $L^{q}$ norms to $L^{p}$ norms. A semigroup $(T_{t})_{t\geq 0}$ is hypercontractive if for all real-valued functions $f$,
$\|T_{t}f\|_{p}\leq\|f\|_{q},$
for $1<q<p<\infty$, whenever $t\geq\tau(p,q)$. These inequalities are deeply connected to logarithmic Sobolev inequalities [Gro75] and underlie results on quantum information theory [Mon12], diffusion processes [BÉ06], and mixing times [BT03].
On the Boolean hypercube $\{0,1\}^{n}$, a canonical example is the Bonami–Beckner semigroup $T_{\rho}$, where $0\leq\rho\leq 1$, defined as
$T_{\rho}f(x)=\mathbb{E}[f(y)],\quad\text{where }y\sim_{\rho}x,$
and $y$ is obtained by flipping each bit of $x$ independently with probability $\frac{1-\rho}{2}$. The operator satisfies the sharp hypercontractive inequality [Bon70, Nel73, Bec75, Gro75]:
$\|T_{\rho}f\|_{p}\leq\|f\|_{q}\quad\text{if }\rho\leq\sqrt{\frac{q-1}{p-1}}$
for $1<q<p<\infty$. This inequality underpins many foundational results in the analysis of Boolean functions, such as the KKL theorem [MOS13], the invariance principle [MOS13].
In contrast, reverse hypercontractivity, introduced by Mossel, Oleszkiewicz, and Sen [MOS13], captures a complementary “anti-smoothing” phenomenon. A semigroup $(T_{t})_{t\geq 0}$ is reverse hypercontractive if for all non-negative functions $f$,
$\|T_{t}f\|_{q}\geq\|f\|_{p},$
for $10<q<p<$ 1, whenever $t\geq\tau(p,q)$. This inequality lower-bounds the dispersion of mass under the semigroup and connects to reverse log-Sobolev inequalities [MOS13], Gaussian isoperimetry [MN15], and tail bounds in correlated settings.
On the Boolean hypercube, the Bonami–Beckner semigroup $T_{\rho}$ also satisfies reverse hypercontractive inequalities [MOS13]: for all non-negative $f:\{0,1\}^{n}\to\mathbb{R}_{\geq 0}$,
$\|T_{\rho}f\|_{q}\geq\|f\|_{p}\quad\text{if }\rho<\sqrt{\frac{1-p}{1-q}}$
for all $10<q<p<$ 1.
Reverse hypercontractivity has enabled several significant applications on the hypercube. It yields anticoncentration and noise stability bounds for Boolean functions of low influence [MOO05], and dimension-free Gaussian isoperimetric inequalities [MN15].
The operator $A_{t}$ serves as the analogue of the Bonami–Beckner operator on the sphere. In this section, we demonstrate that $A_{t}$ is close to the Poisson Markov semigroup $P_{t}$ in the $L^{2}$ norm. Using the logarithmic Sobolev inequality for $P_{t}$, we then establish that $P_{t}$ satisfies reverse hypercontractivity inequalities, which in turn imply reverse hypercontractivity inequalities for $A_{t}$.
3.1 Poisson Markov Semigroup
Definition 3.1.
A family $(P_{t})_{t\geq 0}$ of operators on real-valued measurable functions on ${\mathbb{S}^{n-1}}$ is called a Markov semigroup if it satisfies the following properties:
  1. (i)
    For each $t\geq 0$, $P_{t}$ is a linear operator mapping bounded measurable functions to bounded measurable functions.
  2. (ii)
    $P_{t}(1)=1$, where $1$ denotes the constant function on ${\mathbb{S}^{n-1}}$ (mass conservation).
  3. (iii)
    If $f\geq 0$, then $P_{t}f\geq 0$ (positivity preservation).
  4. (iv)
    $P_{0}=\mathrm{Id}$, the identity operator.
  5. (v)
    For any $s,t\geq 0$, $P_{s+t}=P_{s}\circ P_{t}$ (semigroup property).
  6. (vi)
    For each $f$, the map $t\mapsto P_{t}f$ is continuous.
Operators satisfying properties (i)–(iii) are called Markov operators. For any function $f\in L^{2}({\mathbb{S}^{n-1}})$, we can express it in terms of the spherical harmonic basis as
$f(x)=\sum_{k,l}\widehat{f_{k,l}}Y_{k,l}(x),$
where $\widehat{f_{k,l}}=\langle f,Y_{k,l}\rangle$. Define the Poisson semigroup [^1] In the literature, e.g., [Bec92], the Poisson semigroup is typically defined by $P_{r}f(x)=\sum_{k,l}r^{k}\widehat{f_{k,l}}Y_{k,l}(x)$ with multiplicative semigroup property $P_{rs}=P_{r}\circ P_{s}$. We adopt a slightly modified definition here to match the additive semigroup convention of Markov semigroups. $(P_{t})_{t\geq 0}:L^{2}({\mathbb{S}^{n-1}})\to L^{2}({\mathbb{S}^{n-1}})$ by
$P_{t}f(x)=\sum_{k,l}e^{-kt}\widehat{f_{k,l}}Y_{k,l}(x).$
As shown in [Bec92], this can equivalently be written in terms of the Poisson kernel
$K_{r}(x,y)=\frac{1-r^{2}}{|x-ry|^{n}},\quad(-1\leq r\leq 1),$
so that
[…]
It is straightforward to verify that $P_{t}$ satisfies the Markov semigroup properties. Define the (unbounded) operator $L$ of the semigroup by
$Lf(x):=-\left.\frac{d}{dt}P_{t}f(x)\right|_{t=0}=\sum_{k,l}k\,\widehat{f_{k,l}}Y_{k,l}(x).$
The operator $L$ is called the generator of the Markov semigroup $P_{t}$. Then, by the semigroup property, for any $f\in L^{2}({\mathbb{S}^{n-1}})$, we have
$\frac{d}{dt}P_{t}f=-LP_{t}f=-P_{t}Lf,$
and the semigroup can be expressed formally as $P_{t}=e^{-tL}$ via the Taylor expansion of the exponential operator.
We now show that $A_{t}$ is close to $P_{t}$ in the $L^{2}$-norm.
Lemma 3.2.
For any $f\in L^{2}({\mathbb{S}^{n-1}})$,
$\|A_{t}f-P_{t}f\|_{2}=O_{t}(n^{-2/3})\|f\|_{2}$
as $n\to\infty$.
Proof.
Expressing $f$ in the spherical harmonic basis as $f(x)=\sum_{k,l}\widehat{f_{k,l}}Y_{k,l}(x)$, we have
$\|A_{t}f-P_{t}f\|_{2}=\left\|\sum_{k,l}(\mu_{k,t}-e^{-kt})\widehat{f_{k,l}}Y_{k,l}\right\|_{2}=\sum_{k,l}(\mu_{k,t}-e^{-kt})^{2}\left(\widehat{f_{k,l}}\right)^{2}.$
By Lemma 2.2, for $k\leq n^{1/3}$,
$(\mu_{k,t}-e^{-kt})^{2}=O_{t}(n^{-2/3}).$
By Lemma 2.3, for $k\geq n^{1/3}$,
$(\mu_{k,t}-e^{-kt})^{2}\leq\left(O\left(\frac{1}{n}\right)+e^{-tn^{1/3}}\right)^{2}=O(n^{-2}).$
Combining both regimes, we conclude
$\|A_{t}f-P_{t}f\|_{2}^{2}\leq O(n^{-2/3})\sum_{k,l}\left(\widehat{f_{k,l}}\right)^{2}=O_{t}(n^{-2/3})\|f\|_{2}^{2}\ .\qed$
3.2 Reverse Hypercontractivity
Mossel, Oleszkiewicz, and Sen [MOS13] showed that for Markov semigroups, reverse hypercontractivity inequalities follow from log-Sobolev inequalities. We will use this fact to establish reverse hypercontractivity for the Poisson semigroup $P_{t}$. For any positive function $f>0$, define its entropy by
$\mathop{\bf Ent\/}(f)=\mathop{\mathbb{E}\/}[f\log f]-\mathop{\mathbb{E}\/}[f]\cdot\log\mathop{\mathbb{E}\/}[f].$
For a Markov semigroup $(P_{t})_{t\geq 0}$ with generator $L$, the Dirichlet form is defined as
$\mathcal{E}(f,g)=\mathop{\mathbb{E}\/}[fLg]=\mathop{\mathbb{E}\/}[gLf]=\mathcal{E}(g,f)=-\left.\frac{d}{dt}\mathop{\mathbb{E}\/}[fP_{t}g]\right|_{t=0}.$
Lemma 3.3 ([Bec92], Theorem 1).
The Poisson semigroup $(P_{t})_{t\geq 0}$ satisfies the log-Sobolev inequality:
$\mathop{\bf Ent\/}(f^{2})\leq C\mathcal{E}(f,f),$
for some constant $C\leq\frac{1}{2}$.
Lemma 3.4 ([MOS13], Theorem 1.10).
If a Markov semigroup $(P_{t})_{t\geq 0}$ satisfies the log-Sobolev inequality with constant $C$, then for all $q<p<1$ and every positive function $f:{\mathbb{S}^{n-1}}\to\mathbb{R}$, the following inequality holds for all $t\geq\frac{C}{4}\log\frac{1-q}{1-p}$:
$\|P_{t}f\|_{q}\geq\|f\|_{p}.$
We now deduce a reverse hypercontractivity inequality for the operator $A_{t}$, leveraging its proximity to $P_{t}$ in the $L^{2}$ norm.
Theorem 3.5.
For any positive functions $f,g\in L^{2}({\mathbb{S}^{n-1}})$, the following reverse hypercontractivity inequality holds:
$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(y)]=\langle f,A_{t}g\rangle\geq\|f\|_{p}\|g\|_{p}-O_{t}(n^{-2/3})\|f\|_{2}\|g\|_{2},$
as $n\to\infty$, for all $0<p\leq 1-e^{-4t}$.
Proof.
By Lemma 3.2, we have
$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(x)]=\langle f,A_{t}g\rangle\geq\langle f,P_{t}g\rangle-O_{t}(n^{-2/3})\|f\|_{2}\|g\|_{2}.$
Let $p^{\prime}=\frac{p}{p-1}$ be the Hölder conjugate of $p$, so that $11/p+1/p^{\prime}=$ 1. Applying the reverse Hölder inequality (see [HLP52], Theorem 13), we obtain
$\langle f,P_{t}g\rangle=\|fP_{t}g\|_{1}\geq\|f\|_{p}\|P_{t}g\|_{p^{\prime}}.$
From Lemma 3.3, the semigroup $(P_{t})_{t\geq 0}$ satisfies a log-Sobolev inequality with constant $C\leq\frac{1}{2}$. Thus, by Lemma 3.4,
$\|P_{t}g\|_{p^{\prime}}\geq\|g\|_{p}$
provided that $t\geq\frac{1}{8}\log\frac{1-p^{\prime}}{1-p}=-\frac{1}{4}\log(1-p)$. Therefore, the inequality holds for all $p\leq 1-e^{-4t}$, completing the proof. ∎
4 Density Sphere Avoidance
In this section, we present our main results: the density version of the Frankl–Rödl theorem on the sphere (Theorem 4.3) and the density sphere avoidance result for inductive configurations (Theorem 4.6).
4.1 Inductive Configurations
Let $(v_{1},\cdots,v_{k})$ be a configuration on ${\mathbb{S}^{n-1}}$, where $v_{i}\in{\mathbb{S}^{n-1}}$ are distinct vectors satisfying $\left\langle{v_{i},v_{j}}\right\rangle=r_{i}$ for $i<j$. Specifically, we consider configurations whose covariance matrix has the form
$R(r_{1},\cdots,r_{k-1}):=(v_{1},\cdots,v_{k})^{T}(v_{1},\cdots,v_{k})=\begin{pmatrix}1&r_{1}&r_{1}&\cdots&r_{1}&r_{1}\\ r_{1}&1&r_{2}&\cdots&r_{2}&r_{2}\\ r_{1}&r_{2}&1&\cdots&r_{3}&r_{3}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ r_{1}&r_{2}&r_{3}&\cdots&1&r_{k-1}\\ r_{1}&r_{2}&r_{3}&\cdots&r_{k-1}&1\end{pmatrix}.$
Let $\Delta(n,R)=\{(x_{1},\cdots,x_{k}):(x_{1},\cdots,x_{k})^{T}(x_{1},\cdots,x_{k})=R\}\subseteq({\mathbb{S}^{n-1}})^{k}$ denote the set of tuples that are congruent to the configuration $(v_{1},\cdots,v_{k})$ defined by $R$. We refer to such configurations as inductive configurations, as they can be constructed recursively as follows:
Start with any vector $v_{1}$ and a choice of $-1<r_{1}<1$. Choose $v_{2}$ on the $(n-2)$-dimensional subsphere $S_{v_{1},r_{1}}=\{x\in{\mathbb{S}^{n-1}}:\left\langle{x,v_{1}}\right\rangle=r_{1}\}$. Next, select $r_{2}$ such that the intersection $S_{v_{1},r_{1}}\cap S_{v_{2},r_{2}}$ is nonempty, and choose $v_{3}$ on this $(n-3)$-dimensional subsphere. Proceed inductively: for each $v_{i}$, choose it on the $(n-i)$-dimensional subsphere $\bigcap_{j<i}S_{v_{j},r_{j}}$.
Throughout the paper, we exclude a special case of inductive configuration–configurations where $v_{k}-v_{k-1}$ is the diameter of $\bigcap_{j<k-1}S_{v_{j},r_{j}}$–for reasons that will become clear in the later proof. Notably, both the length of $v_{k}-v_{k-1}$ and the diameter of $\bigcap_{j<k-1}S_{v_{j},r_{j}}$ are independent of the ambient dimension $n$, so this is a constraint on the configuration $R$ itself.
We will show that, fixing any inductive configuration $R$ (excluding the special case), any set $A\subseteq{\mathbb{S}^{n-1}}$ of constant spherical measure will, as $n\to\infty$, contain many congruent copies of $R$. Moreover, we provide an explicit lower bound on the probability that a randomly chosen congruent copy of $R$ lies entirely within $A$.
4.2 Configuration with Pairwise Orthogonal Vectors
Let $\Delta_{k}(n,0):=\Delta(n,I_{k})=\{(x_{1},\cdots,x_{k})\in({\mathbb{S}^{n-1}})^{k}:\left\langle{v_{i},v_{j}}\right\rangle=0,\ \forall i\neq j\}$ denote the set of all $k$-tuples of pairwise orthogonal vectors in ${\mathbb{S}^{n-1}}$.
Lemma 4.1 ([RK11], Theorem 2.2).
For any measurable set $A\subseteq{\mathbb{S}^{n-1}}$, the condition
$\left|\frac{\sigma_{x,0}(A\cap S_{x,0})}{\sigma(A)}\right|\leq 0.1$
holds for all $x\in{\mathbb{S}^{n-1}}$, except on a subset of measure at most $C\exp(-cn^{1/3})$, where $C,c>0$ are universal constants.
Theorem 4.2.
Fix an integer $k\geq 2$. For any measurable set $A\subseteq{\mathbb{S}^{n-1}}$, the probability that a tuple $(v_{1},\cdots,v_{k})$ drawn uniformly at random from $\Delta_{k}(n,0)$ lies entirely in $A$ is at least
$\mathop{\bf Pr\/}_{(x_{1},\cdots,x_{k})\in\Delta_{k}(n,0)}(x_{1}\in A,\cdots,x_{k}\in A)\geq\Omega\left(\sigma(A)^{k}-\exp(-cn^{1/3})\sigma(A)^{k-1}\right),$
where $c>0$ is a universal constant.
Proof.
We proceed by induction on $k$. Let $B\subseteq{\mathbb{S}^{n-1}}$ denote the set of all $x_{1}$ for which
$\left|\frac{\sigma_{x_{1},0}(A\cap S_{x_{1},0})}{\sigma(A)}\right|\leq 0.1.$
By Lemma 4.1, we have $\sigma(B)\geq 1-C\exp(-cn^{1/3})$, where $C,c>0$ are universal constants.
When $k=2$,
$\mathop{\bf Pr\/}_{(x_{1},x_{2})\in\Delta_{2}(n,0)}(x_{1}\in A,x_{2}\in A)$
[…]
When $k>2$, assume the claim holds for $k-1$. Then:
$\mathop{\bf Pr\/}_{(x_{1},\cdots,x_{k})\in\Delta_{k}(n,0)}(x_{1}\in A,\cdots,x_{k}\in A)$
[…]
By the inductive hypothesis applied within the subsphere $S_{x_{1},0}$, we have:
$\mathop{\bf Pr\/}(x_{2},\cdots,x_{k}\in A\cap S_{x_{1},0}\mid x_{1}\in A\cap B)$
[…]
Multiplying with $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\geq\sigma(A)-C\exp(-cn^{1/3})$, we obtain:
$\mathop{\bf Pr\/}(x_{1},\cdots,x_{k}\in A)$
[…]
completing the induction. ∎
4.3 Main Results
Theorem 4.3.
For any measurable sets $A,B\subseteq{\mathbb{S}^{n-1}}$ and any $r\in(-1,1)$, we have
$\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,\ y\in B)\geq(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O(n^{-2/3})\sqrt{\sigma(A)\sigma(B)}.$
Proof.
The case $r=0$ follows from a stronger version of Lemma 4.1, given by [RK11, Theorem 5.1]. Now consider $r\neq 0$. Let
$f=\mathbbm{1}_{A},\quad\text{and}\quad g(x)=\begin{cases}\mathbbm{1}_{B}(x)&\text{if }r>0,\\ \mathbbm{1}_{B}(-x)&\text{if }r<0.\end{cases}$
Applying Theorem 3.5 with $p=1-r^{4}$, we obtain
$\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,\ y\in B)$ $=\mathop{\mathbb{E}\/}_{x\cdot y=|r|}[f(x)g(y)]$
$\geq\|f\|_{p}\|g\|_{p}-O(n^{-2/3})\|f\|_{2}\|g\|_{2}$
$=(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O(n^{-2/3})\sqrt{\sigma(A)\sigma(B)}\ .\qed$
Fix an inductive configuration $R$ and let $A\subseteq{\mathbb{S}^{n-1}}$ be measurable. For any $x\in{\mathbb{S}^{n-1}}$, recall that $\sigma_{x,r}$ denotes the uniform measure on $S_{x,r}$. We say that a vector $x\in A$ is good if
$\sigma_{x,r}(A\cap S_{x,r})\geq\frac{1}{2}\sigma(A)^{(1+r^{4})/(1-r^{4})}.$
Using Theorem 4.3, we can show that many such good vectors must exist:
Proposition 4.4.
Define the set of good vectors:
$A_{\mathrm{good}}=\left\{x\in{\mathbb{S}^{n-1}}:\sigma_{x,r}(A\cap S_{x,r})\geq\frac{1}{2}\sigma(A)^{(1+r^{4})/(1-r^{4})}\right\}.$
Then
$\sigma(A_{\mathrm{good}})\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}(n^{-2/3})\sigma(A).$
Proof.
By Theorem 4.3, we have
$\mathop{\mathbb{E}\/}_{x}[\sigma_{x,r}(A\cap S_{x,r})\mid x\in A]=\mathop{\bf Pr\/}_{x\cdot y=r}(y\in A\mid x\in A)\geq\sigma(A)^{(1+r^{4})/(1-r^{4})}-O_{r}(n^{-2/3}).$
Split the expectation by conditioning on $A_{\mathrm{good}}$ and its complement:
$\mathop{\mathbb{E}\/}_{x}[\sigma_{x,r}(A\cap S_{x,r})\mid x\in A]$ $=\mathop{\mathbb{E}\/}[\sigma_{x,r}(A\cap S_{x,r})\mid x\in A_{\mathrm{good}}]\mathop{\bf Pr\/}(x\in A_{\mathrm{good}}\mid x\in A)$
$\quad+\mathop{\mathbb{E}\/}[\sigma_{x,r}(A\cap S_{x,r})\mid x\notin A_{\mathrm{good}}]\mathop{\bf Pr\/}(x\notin A_{\mathrm{good}}\mid x\in A)$
$\leq 1\cdot\mathop{\bf Pr\/}(x\in A_{\mathrm{good}}\mid x\in A)+\frac{1}{2}\sigma(A)^{(1+r^{4})/(1-r^{4})}.$
Rearranging yields
$\mathop{\bf Pr\/}(x\in A_{\mathrm{good}}\mid x\in A)\geq\mathop{\mathbb{E}\/}_{x}[\sigma_{x,r}(A\cap S_{x,r})\mid x\in A]-\frac{1}{2}\sigma(A)^{(1+r^{4})/(1-r^{4})},$
and thus
$\sigma(A_{\mathrm{good}})=\mathop{\bf Pr\/}(x\in A_{\mathrm{good}}\mid x\in A)\cdot\sigma(A)\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}(n^{-2/3})\sigma(A).$
The following proposition shows the relation between inner product on ${\mathbb{S}^{n-1}}$ and the normalized inner product on $S_{x,c}$:
Proposition 4.5.
Let $x_{1},x_{2},x_{3}\in{\mathbb{S}^{n-1}}$ be such that $\left\langle{x_{1},x_{2}}\right\rangle=\left\langle{x_{1},x_{3}}\right\rangle=c$ and $\left\langle{x_{2},x_{3}}\right\rangle=r$. For $i=2,3$, we can write
$x_{i}=cx_{1}+\sqrt{1-c^{2}}y_{i},$
where $y_{i}=\frac{x_{i}-cx_{1}}{\|x_{i}-cx_{1}\|_{2}}$ satisfies $\left\langle{y_{i},x_{1}}\right\rangle=0$ and
$\left\langle{y_{2},y_{3}}\right\rangle=f_{c}(r),$
with $f_{c}(r)=\frac{r-c^{2}}{1-c^{2}}$.
Proof.
A direct computation shows that $y_{i}$ is orthogonal to $x_{1}$ and that $\|x_{i}-cx_{1}\|_{2}=\sqrt{1-c^{2}}$. Thus,
$\left\langle{y_{2},y_{3}}\right\rangle$ $=\frac{\left\langle{x_{2}-cx_{1},x_{3}-cx_{1}}\right\rangle}{1-c^{2}}$
$=\frac{\left\langle{x_{2},x_{3}}\right\rangle-c\left\langle{x_{1},x_{3}}\right\rangle-c\left\langle{x_{1},x_{2}}\right\rangle+c^{2}\|x_{1}\|_{2}^{2}}{1-c^{2}}$
$=\frac{r-c^{2}}{1-c^{2}}\ .\qed$
We now apply the above propositions to prove our main result using induction.
Theorem 4.6.
Let $(v_{1},\dots,v_{k})$ be an inductive configuration with covariance matrix $R$, and assume that
$v_{k}-v_{k-1}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right).$
Then for any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega(n^{-2/3})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta(n,R)$ lies entirely in $A$ satisfies
$\mathop{\bf Pr\/}_{(x_{1},\dots,x_{k})\in\Delta(n,R)}(x_{1}\in A,\dots,x_{k}\in A)\geq\Omega_{R}(\sigma(A)^{C_{R}}),$
as $n\to\infty$. The constant $C_{R}$ is given by
$C_{R}=\sum_{i=1}^{k-1}\frac{2}{1-c_{i}^{4}}\prod_{j=1}^{i-1}\frac{1+c_{j}^{4}}{1-c_{j}^{4}},$
with
$c_{1}=r_{1},\quad c_{i}=(f_{c_{i-1}}\circ\cdots\circ f_{c_{1}})(r_{i}),\quad i=2,\dots,k-1,$
and $f_{c}(r)=\frac{r-c^{2}}{1-c^{2}}$.
Proof.
We proceed by induction on $k$. The base case $k=2$ follows directly from Theorem 4.3. The condition $v_{2}-v_{1}\neq\operatorname{diam}({\mathbb{S}^{n-1}})$ ensures $r=\left\langle{v_{1},v_{2}}\right\rangle>-1$.
Assume the result holds for all configurations of size $<k$, and let $c_{1}=r_{1}$. Define
$A_{\mathrm{good}}=\left\{x\in{\mathbb{S}^{n-1}}:\sigma_{x,c_{1}}(A\cap S_{x,c_{1}})\geq\frac{1}{2}\sigma(A)^{(1+c_{1}^{4})/(1-c_{1}^{4})}\right\}.$
Then,
$\mathop{\bf Pr\/}_{(x_{1},\dots,x_{k})\in\Delta(n,R)}(x_{1},\dots,x_{k}\in A)$ $=\mathop{\bf Pr\/}(x_{1}\in A)\cdot\mathop{\bf Pr\/}(x_{2},\dots,x_{k}\in A\mid x_{1}\in A)$
$\geq\mathop{\bf Pr\/}(x_{1}\in A_{\mathrm{good}})\cdot\mathop{\bf Pr\/}(x_{2},\dots,x_{k}\in A\mid x_{1}\in A_{\mathrm{good}}).$
Given $x_{1}$, we write $x_{i}=c_{1}x_{1}+\sqrt{1-c_{1}^{2}}y_{i}$ for $i\geq 2$, where $y_{i}=\frac{x_{i}-r_{1}x_{1}}{\left\lVert x_{i}-r_{1}x_{1}\right\rVert_{2}}$. Conditioned on $x_{1}$, according to Proposition 4.5, the induced configuration $(y_{2},\dots,y_{k})$ follows from a uniform distribution over $\Delta(n-1,R^{\prime})$, where
$R^{\prime}=R(f_{c_{1}}(r_{2}),\dots,f_{c_{1}}(r_{k})).$
Given $x_{1}$, the event $x_{2}\in A,\cdots,x_{k}\in A$ is equal to $y_{2}\in A\cap S_{x_{1},c_{1}},\cdots,y_{k}\in A\cap S_{x_{1},c_{1}}$. Conditioned on $x_{1}\in A_{\mathrm{good}}$, we have $\sigma_{x_{1},c_{1}}(A\cap S_{x_{1},c_{1}})\geq\frac{1}{2}\sigma(A)^{(1+c_{1}^{4})/(1-c_{1}^{4})}$, so by the inductive hypothesis:
$\mathop{\bf Pr\/}(y_{2},\dots,y_{k}\in A\cap S_{x_{1},c_{1}})\geq\Omega_{R}\left(\sigma(A)^{C_{R^{\prime}}(1+c_{1}^{4})/(1-c_{1}^{4})}\right).$
with
$C_{R^{\prime}}=\sum_{i=2}^{k-1}\frac{2}{1-c_{i}^{4}}\prod_{j=2}^{i-1}{\frac{1+c_{j}^{4}}{1-c_{j}^{4}}}$
where $c_{2}=f_{c_{1}}(r_{2})$ and
$c_{i}=(f_{c_{i-1}}\circ\cdots\circ f_{c_{2}})(f_{c_{1}}(r_{i}))=(f_{c_{i-1}}\circ\cdots\circ f_{c_{1}})(r_{i}),\ i=3,\cdots,k-1$
By Proposition 4.4, we also have
$\mathop{\bf Pr\/}(x_{1}\in A_{\mathrm{good}})\geq\Omega_{R}(\sigma(A)^{2/(1-c_{1}^{4})}).$
Multiplying the two bounds gives
$\mathop{\bf Pr\/}(x_{1},\dots,x_{k}\in A)\geq\Omega_{R}\left(\sigma(A)^{C_{R}}\right),$
where
$C_{R}=\frac{2}{1-c_{1}^{4}}+\frac{1+c_{1}^{4}}{1-c_{1}^{4}}C_{R^{\prime}}=\sum_{i=1}^{k-1}\frac{2}{1-c_{i}^{4}}\prod_{j=1}^{i-1}\frac{1+c_{j}^{4}}{1-c_{j}^{4}}\ .\qed$
Remark 4.7.
The condition
$v_{k}-v_{k-1}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right)$
is equivalent to $c_{k-1}\neq-1$, which is necessary for applying Theorem 4.3. For a given configuration $(v_{1},\dots,v_{k})$, the quantity $c_{i}$ corresponds to the inner product between $v_{i}$ and later vectors $v_{j}$, after projecting to the $(n-i)$-subsphere $\bigcap_{l<i}S_{v_{l},r_{l}}$ and normalizing. Since $v_{1},\dots,v_{k}$ are distinct, we have $-1<c_{i}<1$ for all $i\leq k-2$ and $c_{k-1}<1$. The only forbidden case is $c_{k-1}=-1$, which occurs precisely when $v_{k}-v_{k-1}$ equals the diameter of the intersection $\bigcap_{j<k-1}S_{v_{j},r_{j}}$.
A configuration is sphere Ramsey if for any $c>0$, any $c$-coloring of the sphere contains a monochromatic congruent copy of the configuration,. Since any $c$-coloring must contain a monochromatic set $A$ with $\sigma(A)\geq 1/c$, an immediate corollary to Theorem 4.6 is that all inductive configurations (excluding the special case) are sphere Ramsey.
Corollary 4.8.
Let $(v_{1},\dots,v_{k})$ be an inductive configuration with covariance matrix $R$, and assume that
$v_{k}-v_{k-1}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right).$
Then $R$ is sphere Ramsey.
A particularly important class of inductive configurations are $k$-simplices:
$\Delta_{k}(n,r):=\Delta(n,R(r,\dots,r))=\{(x_{1},\dots,x_{k})\in({\mathbb{S}^{n-1}})^{k}:\left\langle{x_{i},x_{j}}\right\rangle=r,\ \forall i\neq j\}.$
In this case, the coefficients $c_{i}$ admit the closed-form expression $c_{i}=\frac{r}{1+(i-1)r}$. The condition $c_{k-1}>-1$ is equivalent to $r>-\frac{1}{k-1}$. We thus obtain the following corollary:
Corollary 4.9.
Fix any $r\in\left(-\frac{1}{k-1},1\right)$. For any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega(n^{-2/3})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)$ lies entirely in $A$ satisfies
$\mathop{\bf Pr\/}_{(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)}(x_{1}\in A,\dots,x_{k}\in A)\geq\Omega_{k,r}(\sigma(A)^{C_{R}}),$
as $n\to\infty$. The constant $C_{R}$ is given by
$C_{R}=\sum_{i=1}^{k-1}\frac{2}{1-c_{i}^{4}}\prod_{j=1}^{i-1}\frac{1+c_{j}^{4}}{1-c_{j}^{4}},$
where
$c_{i}=\frac{r}{1+(i-1)r},\quad\text{for }i=1,\dots,k-1.$
5 Open Problems
Our density version of the Frankl–Rödl theorem on the sphere (Theorem 4.3) requires the set $A\subseteq\mathbb{S}^{n-1}$ to have spherical measure $\sigma(A)>\omega(n^{-2/3})$ in order to guarantee a non-trivial lower bound. In contrast, the classical Frankl–Rödl theorem asserts that for any $-1<r<1$, there exists a constant $\epsilon=\epsilon(r)>0$ such that
$\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,y\in A)>0$
holds for all sets $A$ with $\sigma(A)>\Omega(\epsilon^{n})$. Notably, when $r=0$, our result in Theorem 4.2 recovers the classical theorem for exponentially small sets. A natural open problem is to extend our techniques and obtain explicit lower bounds on
$\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,y\in A)$
for all $-1<r<1$, even when $\sigma(A)$ is exponentially small in $n$.
Another direction concerns the class of configurations we consider. Our density result applies to inductive configurations, but Matoušek and Rödl [MR95] showed that for any configuration $P$ with circumradius less than 1, the probability
$\mathop{\bf Pr\/}_{(x_{1},\ldots,x_{k})\in P}(x_{i}\in A\ \forall i)>0$
is positive for any set $A\subseteq\mathbb{S}^{n-1}$ with constant density $\sigma(A)>\epsilon(P)$. It remains open whether one can establish a density version of this theorem—that is, to derive quantitative lower bounds on the probability in terms of $\sigma(A)$.
Finally, our current results can be almost extended to 3-point configurations. For example, using Theorem 4.3, we can obtain bounds on
$\mathop{\bf Pr\/}_{\begin{subarray}{c}x\cdot y=r_{1}\\ x\cdot z=r_{2}\end{subarray}}(x\in A,y\in A,z\in A),$
but this does not yet yield control over 3-point configurations, which require bounding
$\mathop{\bf Pr\/}_{\begin{subarray}{c}x\cdot y=r_{1}\\ x\cdot z=r_{2}\\ y\cdot z=r_{3}\end{subarray}}(x\in A,y\in A,z\in A).$
It is an open question whether a density theorem can be proved for general 3-point configurations on the sphere.

### v3

@@
 Additionally, we prove a density version of spherical avoidance problems, which generalize from pairwise avoidance to broader configurations with prescribed pairwise inner products.
-Our framework encompasses a class of configurations we call inductive configurations, which include simplices with any prescribed inner product $-1<r<1$.
+Our framework encompasses a class of configurations we call inductive configurations, which include simplices with any prescribed inner product 1<r<1-1<r<1.
 As a consequence of our density statement, we show that all inductive configurations are sphere Ramsey.
@@
 The Frankl–Rödl theorem [FR87] is a foundational result in extremal combinatorics and theoretical computer science.
-It states that for any fixed $10<\gamma<$ 1, assuming $(1-\gamma)n$ is even, if a set $A\subseteq\{-1,1\}^{n}$ contains no pair of distinct points at Hamming distance exactly $(1-\gamma)n$, then the fractional size of $A$ must be exponentially small.
+It states that for any fixed $0<\gamma<1$, assuming $(1-\gamma)n$ is even, if a set $A\subseteq\{-1,1\}^{n}$ contains no pair of distinct points at Hamming distance exactly $(1-\gamma)n$, then the fractional size of $A$ must be exponentially small.
 That is, there exists a constant $\epsilon=\epsilon(\gamma)<1$ such that
@@
 This theorem has found broad applications in theoretical computer science, particularly in the analysis of hardness of approximation.
-The graph ${\mathrm{FR}_{\gamma}^{n}}$ has been used to construct integrality gap instances for problems such as 3-Coloring [KG98, KMS98, Cha02, AG11, KOTZ14] and Vertex-Cover [KG98, Cha02, ABL02, ABLT06, GMT08, GMPT10, KOTZ14].
+The graph $\mathrm{FR}_{\gamma}^{n}$} has been used to construct integrality gap instances for problems such as 3-Coloring [KG98, KMS98, Cha02, AG11, KOTZ14] and Vertex-Cover [KG98, Cha02, ABLT06, GMT08, GMPT10, KOTZ14].
 The Frankl–Rödl theorem extends naturally from pairs of points to general forbidden configurations.
@@
 it is natural to ask whether one can give a quantitative lower bound on this probability in terms of the density $|A|/2^{n}$.
-Benabbas, Hatami, and Magen [BHM12] answered this by proving a density version of the Frankl–Rödl theorem: for any $10<\alpha<$ 1, if $|A|/2^{n}\geq\alpha$, then
+Benabbas, Hatami, and Magen [BHM12] answered this by proving a density version of the Frankl–Rödl theorem: for any $0<\alpha<1$, if $|A|/2^{n}\geq\alpha$, then
 $\mathop{\bf Pr\/}_{\begin{subarray}{c}x,y\in\{-1,1\}^{n}\\ x\cdot y=(2\gamma-1)n\end{subarray}}(x\in A,y\in A)\geq 2\left(\alpha/2\right)^{\frac{2}{1-|2\gamma-1|}}-o(1).$
@@
 al [KOTZ14] showed that the proof can be conducted in the sum-of-squares (SOS) proof system.
-In particular, they showed that for any $40<\gamma\leq 1/$ 4, the SOS/Lasserre SDP hierarchy at degree $4\left\lceil\frac{1}{4\gamma}\right\rceil$ certifies that the maximum independent set in ${\mathrm{FR}_{\gamma}^{n}}$ has fractional size $o(1)$.
-This implies that a degree-4 algorithm from the SOS hierarchy can certify that the ${\mathrm{FR}_{\gamma}^{n}}$ SDP integrality gap instances for 3-Coloring have chromatic number $\sigma(1)$, and that a degree-$\left\lceil 1/\gamma\right\rceil$ SOS algorithm can certify the ${\mathrm{FR}_{\gamma}^{n}}$ SDP integrality gap instances for Min-Vertex-Cover has minimum vertex cover $>1-o(1)$.
+In particular, they showed that for any $0<\gamma\leq 1/4$, the SOS/Lasserre SDP hierarchy at degree $4\left\lceil\frac{1}{4\gamma}\right\rceil$ certifies that the maximum independent set in $\mathrm{FR}_{\gamma}^{n}$} has fractional size $o(1)$.
+This implies that a degree-4 algorithm from the SOS hierarchy can certify that the $\mathrm{FR}_{\gamma}^{n}$} SDP integrality gap instances for 3-Coloring have chromatic number $\omega(1)$, and that a degree-$\left\lceil 1/\gamma\right\rceil$ SOS algorithm can certify the $\mathrm{FR}_{\gamma}^{n}$} SDP integrality gap instances for Min-Vertex-Cover has minimum vertex cover >1o(1)>1-o(1).
 In the continuous setting, the sphere $\mathbb{S}^{n-1}$ provides a natural high-dimensional geometric analogue.
@@
 A configuration is sphere Ramsey if any $c$-coloring of of the sphere, there exists a monochromatic congruent copy of the configuration.
-Matoušek and Rödl [MR95] showed that any configuration $P$ with circumradius less than 1 is sphere Ramsey.
+Castro-Silva et al.’s result implied that all simplectic configurations $\Delta_{k}(n,r)$ with $r>0,\ k\geq 2$ are sphere Ramsey.
+On the other hand, Matoušek and Rödl [MR95] showed that any configuration $P$ with circumradius less than 1 is sphere Ramsey.
 This sphere Ramsey theorem was recently used by Brakensiek, Guruswami, and Sandeep [BGS23] to demonstrate integrality gaps for the basic SDP relaxation of certain promise CSPs: namely Boolean symmetric promise CSPs defined by a single predicate pair that lack Majority or Alternate-Threshold (AT) polymorphisms of all odd arities.
@@
 in terms of $\sigma(A)$. While we do not directly resolve the exact configuration posed by BGS, our result applies to a broad class of configurations that includes closely related structures.
-In this paper, we show that for any $-\frac{1}{k-1}<r<1$, there exists a constant $C=C(r)$ such that
+In this paper, we show the following result:
+Theorem 1.1.
+Fix $k,r$ such that $\frac{1}{k-1}<r<1$, there exists constant $C=C(k,r),\ \epsilon=\epsilon(k,r)$ such that
 $\mathop{\bf Pr\/}_{x_{1},\dots,x_{k}\in\Delta_{k}(n,r)}(x_{i}\in A\ \forall i)\geq\Omega_{k,r}\left(\sigma(A)^{C}\right)$
-for all measurable $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega(n^{-2/3})$.
+for all measurable $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega_{k,r}(n^{-\epsilon})$.
 Our result generalizes the work of Castro-Silva et al.
-to simplices with any $-1<r<1$, and more broadly, to a class of inductive configurations defined in Section 4.
-As a consequence, we show that all inductive configurations are sphere Ramsey.
+to simplices with any $\frac{1}{k-1}<r<1$, and more broadly, to a class of inductive configurations defined in Section 4 (see Theorem 4.6).
+In particular, we show that all inductive configurations are sphere Ramsey.
 Inspired by the techniques of Benabbas et al.
@@
 In Section 2, we analyze the eigen-decomposition of $A_{t}$; in Section 3, we relate $A_{t}$ to the Poisson Markov semigroup $P_{t}$ and establish reverse hypercontractive inequalities.
-Finally, in Section 4, we prove our main results on density versions of the Frankl–Rödl theorem on the sphere and density spherical for inductive configurations.
+Finally, in Section 4, we prove our main results on density versions of the Frankl–Rödl theorem on the sphere and density spherical Ramsey statements for inductive configurations.
 2 Eigen-decomposition via Spherical Harmonics
@@
 Given $p>0$ and a measurable function $f:\mathbb{S}^{n-1}\to\mathbb{R}$, we define its $L^{p}$ norm by
-[…]
-[…]
+$\|f\|_{L^{p}(\mathbb{S}^{n-1})}:=\left(\int_{\mathbb{S}^{n-1}}|f(x)|^{p}\,d\sigma(x)\right)^{1/p},$
+where $\sigma$ denotes the uniform surface measure on $\mathbb{S}^{n-1}$, normalized such that $\int_{\mathbb{S}^{n-1}}d\sigma(x)=1$.
+It follows that for $p\geq q>0$, we have $L^{p}(\mathbb{S}^{n-1})\subseteq L^{q}(\mathbb{S}^{n-1})$.
 We focus in particular on $L^{2}(\mathbb{S}^{n-1})$, the space of square-integrable functions.
-It admits an orthonormal basis consisting of spherical harmonics $\{S_{k,\ell}\}$, indexed by integers $k\geq 0$ (the degree) and $1\leq\ell\leq c_{k,n}$, where
+It admits an orthonormal basis consisting of spherical harmonics $\{Y_{k,\ell}\}$, indexed by integers $k\geq 0$ (the degree) and $1\leq\ell\leq c_{k,n}$, where
 $c_{k,n}=\binom{k+n-2}{k}+\binom{k+n-3}{k-1}$.
-Each $S_{k,\ell}$ is the restriction to $\mathbb{S}^{n-1}$ of a harmonic, homogeneous polynomial of degree $k$ on $\mathbb{R}^{n}$.
+Each $Y_{k,\ell}$ is the restriction to $\mathbb{S}^{n-1}$ of a harmonic, homogeneous polynomial of degree $k$ on $\mathbb{R}^{n}$.
 We denote by $\mathcal{S}_{k}\subseteq L^{2}(\mathbb{S}^{n-1})$ the finite-dimensional subspace spanned by spherical harmonics of degree $k$:
-$\mathcal{S}_{k}:=\operatorname{span}\{S_{k,1},\dotsc,S_{k,c_{k,n}}\}.$
+$\mathcal{S}_{k}:=\operatorname{span}\{Y_{k,1},\dotsc,Y_{k,c_{k,n}}\}.$
 Notably, $\mathcal{S}_{0}$ consists of the constant functions. Rotations $U\in\mathrm{SO}(n)$ act on $f\in L^{2}(\mathbb{S}^{n-1})$ via
@@
 For $t\geq 0$, define an operator $A_{t}:L^{2}(\mathbb{S}^{n-1})\to L^{2}(\mathbb{S}^{n-1})$ by
-[…]
-where $\sigma_{x,r}$ is the uniform probability measure on the $(n-2)$-subsphere $\{y\in\mathbb{S}^{n-1}:x\cdot y=r\}$.
-The operator $A_{t}$ is easily seen to commute with rotations. It is also self-adjoint:
+$(A_{t}f)(x):=\int_{\mathbb{S}^{n-1}}f(y)\,d\sigma_{x,e^{-t}}(y)=\mathop{\mathbb{E}\/}_{\begin{subarray}{c}y\in{\mathbb{S}^{n-1}}\\ \ x\cdot y=e^{-t}\end{subarray}}[f(y)],$
+where $\sigma_{x,r}$ is the uniform probability measure on the (n2)(n-2)-subsphere $S_{x,r}:=\{y\in\mathbb{S}^{n-1}:x\cdot y=r\}$.
+The operator $A_{t}$ is easily seen to commute with rotations via
+$(A_{t}Uf)(x)=\mathop{\mathbb{E}\/}_{\begin{subarray}{c}y\in{\mathbb{S}^{n-1}}\\ \ x\cdot y=e^{-t}\end{subarray}}[f(U^{-1}y)]=\mathop{\mathbb{E}\/}_{\begin{subarray}{c}z\in{\mathbb{S}^{n-1}}\\ \ x\cdot Uz=e^{-t}\end{subarray}}[f(z)]=\mathop{\mathbb{E}\/}_{\begin{subarray}{c}z\in{\mathbb{S}^{n-1}}\\ \ U^{-1}x\cdot z=e^{-t}\end{subarray}}[f(z)]=(UA_{t}f)(x)$
+for any $U\in\mathrm{SO}(n)$. It is also self-adjoint by
 $\left\langle{A_{t}f,g}\right\rangle=\mathop{\mathbb{E}\/}_{x\in{\mathbb{S}^{n-1}}}[g(x)A_{t}f(x)]=\mathop{\mathbb{E}\/}_{x\in{\mathbb{S}^{n-1}}}\left[\mathop{\mathbb{E}\/}_{y:x\cdot y=e^{-t}}[g(x)f(y)]\right]=\mathop{\mathbb{E}\/}_{y\in{\mathbb{S}^{n-1}}}\left[\mathop{\mathbb{E}\/}_{x:x\cdot y=e^{-t}}[g(x)f(y)]\right]=\left\langle{f,A_{t}g}\right\rangle.$
@@
 Proof.
-By Schur’s Lemma, any linear operator commuting with the group action must act as a scalar on each irreducible representation.
+By Schur’s Lemma (see e.g.
+[S+77]), any linear operator commuting with the group action must act as a scalar on each irreducible representation.
 Since $A_{t}$ commutes with rotations and $\mathcal{S}_{k}$ are inequivalent irreducible representations for different $k$, the restriction of $A_{t}$ to $\mathcal{S}_{k}$ must be a scalar multiple of the identity.
@@
 2.2 Eigenvalue Estimates
-We now estimate the eigenvalues $\mu_{k,t}$ for different regimes of $k$:
+We now estimate the eigenvalues $\mu_{k,t}$.
 Lemma 2.2.
-For all $0\leq k\leq n^{1/3}$, as $n\to\infty$, the eigenvalues satisfy
-$|\mu_{k,t}-e^{-kt}|=O_{t}(n^{-1/3}),$
-where the implicit constant depends only on $t$ but is independent of $k$.
+The eigenvalues $\mu_{k,t}$ satisfy
+$|\mu_{k,t}-e^{-kt}|=O_{t}\left(\frac{1}{n}\right),$
+as $n\to\infty$, where the implicit constant depends only on $t$ but is independent of $k$.
 Proof.
 Let us first analyze the $k$-th moments of $X_{1}$.
-Notice that the density of $X_{1}$ is supported in $[-1,1]$ and proportional to $(1-x^{2})^{(n-4)/2}$.
+Notice that the density of $X_{1}$ is supported in [1,1][-1,1] and proportional to $(1-x^{2})^{(n-4)/2}$.
 Since the density $X_{1}$ is an odd function, the odd moments vanishes, $\mathop{\mathbb{E}\/}[X_{1}^{k}]=0$ for odd $k$.
@@
 $\mathop{\mathbb{E}\/}[X_{1}^{k}]\leq\left(\frac{k}{n-4}\right)^{k/2}$
-given by [RK11] Lemma 5.5. A direct computation gives
+given by [RK11] Lemma 5.5. We will estimate the eigenvalues $\mu_{k,t}$ for different regimes of $k$.
+When $k\leq\frac{n}{10}$, a direct computation gives
 $\mu_{k,t}$ $=G_{k}(e^{-t})$
@@
 $|\mu_{k,t}-e^{-kt}|\leq{}$ $\sum_{a=1}^{\left\lfloor k/2\right\rfloor}(1-e^{-2t})^{a}e^{-t(k-2a)}\left(\frac{2a}{n-4}\right)^{a}$
-[…]
-[…]
-[…]
-[…]
-[…]
-Lemma 2.3.
-For all $k\geq n^{1/3}$, as $n\to\infty$, the eigenvalues satisfy
-$|\mu_{k,t}|=O\left(\frac{1}{n}\right).$
-Proof.
-Since $\mathop{\mathbb{E}\/}[X_{1}^{2}]\leq\frac{2}{n-4}$, by Markov’s inequality,
-$\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})=\mathop{\bf Pr\/}(|X_{1}|^{2}\geq\frac{1}{4})\leq\frac{8}{n-4},$
+$\leq{}$ $\frac{2}{n-4}+\sum_{a=2}^{\left\lfloor k/2\right\rfloor}(1-e^{-2t})^{a}e^{-t(k-2a)}\left(\frac{2a}{n-4}\right)^{a}$
+$\leq{}$ $\frac{2}{n-4}+\sum_{a=2}^{\left\lfloor k/2\right\rfloor}\left(\frac{2a}{n-4}\right)^{a}$
+Since $\left(\frac{2a}{n-4}\right)^{a}$ is decreasing with $a$ when $2\leq a\leq\left\lfloor k/2\right\rfloor\leq\frac{n}{20}$, we can bound
+$|\mu_{k,t}-e^{-kt}|\leq\frac{2}{n-4}+(\left\lfloor k/2\right\rfloor-2)\left(\frac{4}{n-4}\right)^{2}\leq\frac{2}{n-4}+\frac{n}{20}\left(\frac{4}{n-4}\right)^{2}\leq O\left(\frac{1}{n}\right).$
+For the case where $k\geq\frac{n}{10}$, by Markov’s inequality
+$\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})\leq\frac{\mathop{\mathbb{E}\/}\left[|X_{1}|^{2\left\lceil n/20\right\rceil}\right]}{(1/2)^{2\left\lceil n/20\right\rceil}}\leq\left(\frac{4\left\lceil n/20\right\rceil}{n-4}\right)^{n/20}\leq O\left(\frac{1}{n}\right)$
 which leads us to
-$|\mu_{k,t}|$ $=\ |G_{k}(e^{-t})|$
-$=\ \left|\mathop{\mathbb{E}\/}\left[\left(e^{-t}+iX_{1}\sqrt{1-e^{-2t}}\right)^{k}\right]\right|$
-$\leq\ \mathop{\mathbb{E}\/}\left[\left|e^{-t}+iX_{1}\sqrt{1-e^{-2t}}\right|^{k}\right]$
-$\leq\ \mathop{\mathbb{E}\/}\left[\left(e^{-2t}+(1-e^{-2t})|X_{1}|\right)^{k}\right]$
-$\leq\ \mathop{\bf Pr\/}(|X_{1}|<\frac{1}{2})\left(e^{-2t}+\frac{1}{2}(1-e^{-2t})\right)^{k}+\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})$
-$=\ O\left(\left(\frac{1+e^{-2t}}{2}\right)^{n^{1/3}}\right)+O(\frac{1}{n})\ =\ O(\frac{1}{n})$
+$|\mu_{k,t}-e^{-kt}|$
+$\leq{}$ $|G_{k}(e^{-t})|+|e^{-kt}|$
+$\leq{}$ $\mathop{\mathbb{E}\/}\left[\left|e^{-t}+iX_{1}\sqrt{1-e^{-2t}}\right|^{k}\right]+|e^{-kt}|$
+$\leq{}$ $\mathop{\mathbb{E}\/}\left[\left(e^{-2t}+(1-e^{-2t})|X_{1}|\right)^{k}\right]+|e^{-kt}|$
+$\leq{}$ $\mathop{\bf Pr\/}(|X_{1}|<\frac{1}{2})\left(e^{-2t}+\frac{1}{2}(1-e^{-2t})\right)^{k}+\mathop{\bf Pr\/}(|X_{1}|\geq\frac{1}{2})+|e^{-\frac{n}{10}t}|$
+$\leq{}$ $\ O\left(\left(\frac{1+e^{-2t}}{2}\right)^{n/10}\right)+O(\frac{1}{n})+O_{t}(\frac{1}{n})$
+$\leq{}$ $O_{t}(\frac{1}{n})$
 as $n\rightarrow\infty$. ∎
@@
 $\|T_{t}f\|_{q}\geq\|f\|_{p},$
-for $10<q<p<$ 1, whenever $t\geq\tau(p,q)$.
+for $0<q<p<1$, whenever $t\geq\tau(p,q)$.
 This inequality lower-bounds the dispersion of mass under the semigroup and connects to reverse log-Sobolev inequalities [MOS13], Gaussian isoperimetry [MN15], and tail bounds in correlated settings.
@@
 $\|T_{\rho}f\|_{q}\geq\|f\|_{p}\quad\text{if }\rho<\sqrt{\frac{1-p}{1-q}}$
-for all $10<q<p<$ 1.
+for all $0<q<p<1$.
 Reverse hypercontractivity has enabled several significant applications on the hypercube.
-It yields anticoncentration and noise stability bounds for Boolean functions of low influence [MOO05], and dimension-free Gaussian isoperimetric inequalities [MN15].
+It was used to prove the “It Ain’t Over Till It’s Over” conjecture [Kho02] from social choice theory [MOO05], and dimension-free Gaussian isoperimetric inequalities [MN15].
 The operator $A_{t}$ serves as the analogue of the Bonami–Beckner operator on the sphere.
@@
 so that
-[…]
-It is straightforward to verify that $P_{t}$ satisfies the Markov semigroup properties. Define the (unbounded) operator $L$ of the semigroup by
-$Lf(x):=-\left.\frac{d}{dt}P_{t}f(x)\right|_{t=0}=\sum_{k,l}k\,\widehat{f_{k,l}}Y_{k,l}(x).$
-The operator $L$ is called the generator of the Markov semigroup $P_{t}$. Then, by the semigroup property, for any $f\in L^{2}({\mathbb{S}^{n-1}})$, we have
-$\frac{d}{dt}P_{t}f=-LP_{t}f=-P_{t}Lf,$
-and the semigroup can be expressed formally as $P_{t}=e^{-tL}$ via the Taylor expansion of the exponential operator.
+$P_{t}f(x)=\int_{{\mathbb{S}^{n-1}}}K_{e^{-t}}(x,y)f(y)\,d\sigma(y).$
+It is straightforward to verify that $P_{t}$ satisfies the Markov semigroup properties.
 We now show that $A_{t}$ is close to $P_{t}$ in the $L^{2}$-norm.
@@
 For any $f\in L^{2}({\mathbb{S}^{n-1}})$,
-$\|A_{t}f-P_{t}f\|_{2}=O_{t}(n^{-2/3})\|f\|_{2}$
+$\|A_{t}f-P_{t}f\|_{2}=O_{t}\left(\frac{1}{n}\right)\|f\|_{2}$
 as $n\to\infty$.
@@
 Expressing $f$ in the spherical harmonic basis as $f(x)=\sum_{k,l}\widehat{f_{k,l}}Y_{k,l}(x)$, we have
-$\|A_{t}f-P_{t}f\|_{2}=\left\|\sum_{k,l}(\mu_{k,t}-e^{-kt})\widehat{f_{k,l}}Y_{k,l}\right\|_{2}=\sum_{k,l}(\mu_{k,t}-e^{-kt})^{2}\left(\widehat{f_{k,l}}\right)^{2}.$
-By Lemma 2.2, for $k\leq n^{1/3}$,
-$(\mu_{k,t}-e^{-kt})^{2}=O_{t}(n^{-2/3}).$
-By Lemma 2.3, for $k\geq n^{1/3}$,
-$(\mu_{k,t}-e^{-kt})^{2}\leq\left(O\left(\frac{1}{n}\right)+e^{-tn^{1/3}}\right)^{2}=O(n^{-2}).$
-Combining both regimes, we conclude
-$\|A_{t}f-P_{t}f\|_{2}^{2}\leq O(n^{-2/3})\sum_{k,l}\left(\widehat{f_{k,l}}\right)^{2}=O_{t}(n^{-2/3})\|f\|_{2}^{2}\ .\qed$
+$\|A_{t}f-P_{t}f\|_{2}^{2}=\left\|\sum_{k,l}(\mu_{k,t}-e^{-kt})\widehat{f_{k,l}}Y_{k,l}\right\|_{2}^{2}=\sum_{k,l}|\mu_{k,t}-e^{-kt}|^{2}\left|\widehat{f_{k,l}}\right|^{2}.$
+By Lemma 2.2,
+$|\mu_{k,t}-e^{-kt}|^{2}\leq O_{t}(n^{-2})$
+where we can conclude
+$\|A_{t}f-P_{t}f\|_{2}^{2}\leq O_{t}\left(n^{-2}\right)\sum_{k,l}\left|\widehat{f_{k,l}}\right|^{2}=O_{t}\left(n^{-2}\right)\|f\|_{2}^{2}\ .\qed$
 3.2 Reverse Hypercontractivity
@@
 For any positive functions $f,g\in L^{2}({\mathbb{S}^{n-1}})$, the following reverse hypercontractivity inequality holds:
-$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(y)]=\langle f,A_{t}g\rangle\geq\|f\|_{p}\|g\|_{p}-O_{t}(n^{-2/3})\|f\|_{2}\|g\|_{2},$
+$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(y)]=\langle f,A_{t}g\rangle\geq\|f\|_{p}\|g\|_{p}-O_{t}\left(\frac{1}{n}\right)\|f\|_{2}\|g\|_{2},$
 as $n\to\infty$, for all $0<p\leq 1-e^{-4t}$.
@@
 By Lemma 3.2, we have
-$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(x)]=\langle f,A_{t}g\rangle\geq\langle f,P_{t}g\rangle-O_{t}(n^{-2/3})\|f\|_{2}\|g\|_{2}.$
-Let $p^{\prime}=\frac{p}{p-1}$ be the Hölder conjugate of $p$, so that $11/p+1/p^{\prime}=$ 1. Applying the reverse Hölder inequality (see [HLP52], Theorem 13), we obtain
+$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(x)]=\langle f,A_{t}g\rangle\geq\langle f,P_{t}g\rangle-O_{t}\left(\frac{1}{n}\right)\|f\|_{2}\|g\|_{2}.$
+Let $p^{\prime}=\frac{p}{p-1}$ be the Hölder conjugate of $p$, so that $1/p+1/p^{\prime}=1$. Applying the reverse Hölder inequality (see [HLP52], Theorem 13), we obtain
 $\langle f,P_{t}g\rangle=\|fP_{t}g\|_{1}\geq\|f\|_{p}\|P_{t}g\|_{p^{\prime}}.$
@@
 Let $(v_{1},\cdots,v_{k})$ be a configuration on $\mathbb{S}^{n-1}$}, where $v_{i}\in{\mathbb{S}^{n-1}}$ are distinct vectors satisfying $\left\langle{v_{i},v_{j}}\right\rangle=r_{i}$ for $i<j$.
-Specifically, we consider configurations whose covariance matrix has the form
+Specifically, writing $v_{i}$ as column vectors, we consider configurations whose covariance matrix has the form
 $R(r_{1},\cdots,r_{k-1}):=(v_{1},\cdots,v_{k})^{T}(v_{1},\cdots,v_{k})=\begin{pmatrix}1&r_{1}&r_{1}&\cdots&r_{1}&r_{1}\\ r_{1}&1&r_{2}&\cdots&r_{2}&r_{2}\\ r_{1}&r_{2}&1&\cdots&r_{3}&r_{3}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ r_{1}&r_{2}&r_{3}&\cdots&1&r_{k-1}\\ r_{1}&r_{2}&r_{3}&\cdots&r_{k-1}&1\end{pmatrix}$.
@@
 We refer to such configurations as inductive configurations, as they can be constructed recursively as follows:
-Start with any vector $v_{1}$ and a choice of $-1<r_{1}<1$.
-Choose $v_{2}$ on the $(n-2)$-dimensional subsphere $S_{v_{1},r_{1}}=\{x\in{\mathbb{S}^{n-1}}:\left\langle{x,v_{1}}\right\rangle=r_{1}\}$.
-Next, select $r_{2}$ such that the intersection $S_{v_{1},r_{1}}\cap S_{v_{2},r_{2}}$ is nonempty, and choose $v_{3}$ on this $(n-3)$-dimensional subsphere.
-Proceed inductively: for each $v_{i}$, choose it on the $(n-i)$-dimensional subsphere $\bigcap_{j<i}S_{v_{j},r_{j}}$.
-Throughout the paper, we exclude a special case of inductive configuration–configurations where $v_{k}-v_{k-1}$ is the diameter of $\bigcap_{j<k-1}S_{v_{j},r_{j}}$–for reasons that will become clear in the later proof.
+Start with any vector $v_{1}$ and a choice of $1<r_{1}<1$.
+Choose $v_{2}$ on the (n2)(n-2)-subsphere $S_{v_{1},r_{1}}=\{x\in{\mathbb{S}^{n-1}}:\left\langle{x,v_{1}}\right\rangle=r_{1}\}$.
+Next, select $r_{2}$ such that the intersection $S_{v_{1},r_{1}}\cap S_{v_{2},r_{2}}$ is nonempty, and choose $v_{3}$ on this (n3)(n-3)-subsphere.
+Proceed inductively: for each $v_{i}$, choose it on the (ni)(n-i)-subsphere $\bigcap_{j<i}S_{v_{j},r_{j}}$.
+Throughout the paper, we exclude a special case of inductive configuration–configurations where the length of $v_{k}-v_{k-1}$ is the diameter of $\bigcap_{j<k-1}S_{v_{j},r_{j}}$—for reasons that will become clear in the later proof.
 Notably, both the length of $v_{k}-v_{k-1}$ and the diameter of $\bigcap_{j<k-1}S_{v_{j},r_{j}}$ are independent of the ambient dimension $n$, so this is a constraint on the configuration $R$ itself.
@@
 4.2 Configuration with Pairwise Orthogonal Vectors
-Let $\Delta_{k}(n,0):=\Delta(n,I_{k})=\{(x_{1},\cdots,x_{k})\in({\mathbb{S}^{n-1}})^{k}:\left\langle{v_{i},v_{j}}\right\rangle=0,\ \forall i\neq j\}$ denote the set of all $k$-tuples of pairwise orthogonal vectors in ${\mathbb{S}^{n-1}}$.
+Let $\Delta_{k}(n,0):=\Delta(n,I_{k})=\{(x_{1},\cdots,x_{k})\in({\mathbb{S}^{n-1}})^{k}:\left\langle{x_{i},x_{j}}\right\rangle=0,\ \forall i\neq j\}$ denote the set of all $k$-tuples of pairwise orthogonal vectors in $\mathbb{S}^{n-1}$}.
+Recall that $S_{x,r}$ denotes the (n2)(n-2)-subsphere $\{y\in\mathbb{S}^{n-1}:x\cdot y=r\}$ and $\sigma_{x,r}$ is the uniform probability measure on $S_{x,r}$.
 Lemma 4.1 ([RK11], Theorem 2.2).
@@
 For any measurable set $A\subseteq{\mathbb{S}^{n-1}}$, the probability that a tuple $(v_{1},\cdots,v_{k})$ drawn uniformly at random from $\Delta_{k}(n,0)$ lies entirely in $A$ is at least
-$\mathop{\bf Pr\/}_{(x_{1},\cdots,x_{k})\in\Delta_{k}(n,0)}(x_{1}\in A,\cdots,x_{k}\in A)\geq\Omega\left(\sigma(A)^{k}-\exp(-cn^{1/3})\sigma(A)^{k-1}\right),$
-where $c>0$ is a universal constant.
+$\mathop{\bf Pr\/}_{(x_{1},\cdots,x_{k})\in\Delta_{k}(n,0)}(x_{1}\in A,\cdots,x_{k}\in A)\geq\Omega_{k}\left(\sigma(A)^{k}-C_{k}\exp(-cn^{1/3})\sigma(A)^{k-1}\right),$
+where $c>0$ is a universal constant and $C_{k}$ is a constant depending only on $k$.
 Proof.
@@
 $\mathop{\bf Pr\/}_{(x_{1},x_{2})\in\Delta_{2}(n,0)}(x_{1}\in A,x_{2}\in A)$
-[…]
-[…]
-[…]
-[…]
-[…]
-When $k>2$, assume the claim holds for $k-1$. Then:
+$\geq{}$ $\mathop{\bf Pr\/}_{(x_{1},x_{2})\in\Delta_{2}(n,0)}(x_{1}\in A\cap B,x_{2}\in A)$
+$={}$ $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\cdot\mathop{\bf Pr\/}(x_{2}\in A\mid x_{1}\in A\cap B)$
+$={}$ $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\cdot\mathop{\mathbb{E}\/}[\sigma_{x_{1},0}(A\cap S_{x_{1},0})\mid x_{1}\in A\cap B]$
+$\geq{}$ $\left(\sigma(A)-C\exp(-cn^{1/3})\right)\cdot 0.9\sigma(A)$
+$\geq{}$ $\Omega\left(\sigma(A)^{2}-C\exp(-cn^{1/3})\sigma(A)\right)$
+When $k>2$, assume the claim holds for k1k-1. Then:
 $\mathop{\bf Pr\/}_{(x_{1},\cdots,x_{k})\in\Delta_{k}(n,0)}(x_{1}\in A,\cdots,x_{k}\in A)$
-[…]
-[…]
+$\geq{}$ $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\cdot\mathop{\bf Pr\/}(x_{2},\cdots,x_{k}\in A\mid x_{1}\in A\cap B)$
+$={}$ $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\cdot\mathop{\bf Pr\/}(x_{2},\cdots,x_{k}\in A\cap S_{x_{1},0}\mid x_{1}\in A\cap B).$
 By the inductive hypothesis applied within the subsphere $S_{x_{1},0}$, we have:
 $\mathop{\bf Pr\/}(x_{2},\cdots,x_{k}\in A\cap S_{x_{1},0}\mid x_{1}\in A\cap B)$
-[…]
-[…]
+$\geq{}$ $\Omega_{k}\left((0.9\sigma(A))^{k-1}-C_{k}\exp(-cn^{1/3})(0.9\sigma(A))^{k-2}\right)$
+$\geq{}$ $\Omega_{k}\left(\sigma(A)^{k-1}-C_{k}\exp(-cn^{1/3})\sigma(A)^{k-2}\right).$
 Multiplying with $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\geq\sigma(A)-C\exp(-cn^{1/3})$, we obtain:
 $\mathop{\bf Pr\/}(x_{1},\cdots,x_{k}\in A)$
-[…]
-[…]
+$\geq{}$ $\Omega_{k}\left(\sigma(A)^{k-1}-C_{k}\exp(-cn^{1/3})\sigma(A)^{k-2}\right)\cdot\left(\sigma(A)-C\exp(-cn^{1/3})\right)$
+$\geq{}$ $\Omega_{k}\left(\sigma(A)^{k}-C^{\prime}_{k}\exp(-cn^{1/3})\sigma(A)^{k-1}\right),$
 completing the induction. ∎
 4.3 Main Results
-Theorem 4.3.
+We first prove a two-set version of the density sphere Frankl-Rödl theorem which will be later used for induction on inductive configurations.
+Theorem 4.3 (Density Sphere Frankl-Rödl).
 For any measurable sets $A,B\subseteq{\mathbb{S}^{n-1}}$ and any $r\in(-1,1)$, we have
-$\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,\ y\in B)\geq(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O(n^{-2/3})\sqrt{\sigma(A)\sigma(B)}.$
+$\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,\ y\in B)\geq(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right)\sqrt{\sigma(A)\sigma(B)}.$
 Proof.
@@
 $\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,\ y\in B)$ $=\mathop{\mathbb{E}\/}_{x\cdot y=|r|}[f(x)g(y)]$
-$\geq\|f\|_{p}\|g\|_{p}-O(n^{-2/3})\|f\|_{2}\|g\|_{2}$
-$=(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O(n^{-2/3})\sqrt{\sigma(A)\sigma(B)}\ .\qed$
+$\geq\|f\|_{p}\|g\|_{p}-O_{r}\left(\frac{1}{n}\right)\|f\|_{2}\|g\|_{2}$
+$=(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right)\sqrt{\sigma(A)\sigma(B)}\ .\qed$
 Fix an inductive configuration $R$ and let $A\subseteq{\mathbb{S}^{n-1}}$ be measurable.
@@
 $\sigma_{x,r}(A\cap S_{x,r})\geq\frac{1}{2}\sigma(A)^{(1+r^{4})/(1-r^{4})}.$
-Using Theorem 4.3, we can show that many such good vectors must exist:
-Proposition 4.4.
 Define the set of good vectors:
 $A_{\mathrm{good}}=\left\{x\in{\mathbb{S}^{n-1}}:\sigma_{x,r}(A\cap S_{x,r})\geq\frac{1}{2}\sigma(A)^{(1+r^{4})/(1-r^{4})}\right\}.$
-Then
-$\sigma(A_{\mathrm{good}})\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}(n^{-2/3})\sigma(A).$
+Using Theorem 4.3, we can show that many such good vectors must exist.
+Proposition 4.4.
+The set of good vectors satisfies
+$\sigma(A_{\mathrm{good}})\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right)\sigma(A).$
 Proof.
 By Theorem 4.3, we have
-$\mathop{\mathbb{E}\/}_{x}[\sigma_{x,r}(A\cap S_{x,r})\mid x\in A]=\mathop{\bf Pr\/}_{x\cdot y=r}(y\in A\mid x\in A)\geq\sigma(A)^{(1+r^{4})/(1-r^{4})}-O_{r}(n^{-2/3}).$
+$\mathop{\mathbb{E}\/}_{x}[\sigma_{x,r}(A\cap S_{x,r})\mid x\in A]=\mathop{\bf Pr\/}_{x\cdot y=r}(y\in A\mid x\in A)\geq\sigma(A)^{(1+r^{4})/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right).$
 Split the expectation by conditioning on $A_{\mathrm{good}}$ and its complement:
@@
 and thus
-$\sigma(A_{\mathrm{good}})=\mathop{\bf Pr\/}(x\in A_{\mathrm{good}}\mid x\in A)\cdot\sigma(A)\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}(n^{-2/3})\sigma(A).$
+$\sigma(A_{\mathrm{good}})=\mathop{\bf Pr\/}(x\in A_{\mathrm{good}}\mid x\in A)\cdot\sigma(A)\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right)\sigma(A)\ .\qed$
 The following proposition shows the relation between inner product on $\mathbb{S}^{n-1}$} and the normalized inner product on $S_{x,c}$:
@@
 Let $(v_{1},\dots,v_{k})$ be an inductive configuration with covariance matrix $R$, and assume that
-$v_{k}-v_{k-1}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right).$
-Then for any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega(n^{-2/3})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta(n,R)$ lies entirely in $A$ satisfies
+$\left\lVert v_{k}-v_{k-1}\right\rVert_{2}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right).$
+Then for any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega_{R}(n^{-\epsilon_{R}})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta(n,R)$ lies entirely in $A$ satisfies
 $\mathop{\bf Pr\/}_{(x_{1},\dots,x_{k})\in\Delta(n,R)}(x_{1}\in A,\dots,x_{k}\in A)\geq\Omega_{R}(\sigma(A)^{C_{R}}),$
@@
 $c_{1}=r_{1},\quad c_{i}=(f_{c_{i-1}}\circ\cdots\circ f_{c_{1}})(r_{i}),\quad i=2,\dots,k-1,$
-and $f_{c}(r)=\frac{r-c^{2}}{1-c^{2}}$.
+and $f_{c}(r)=\frac{r-c^{2}}{1-c^{2}}$. The constant $\epsilon_{R}$ is given by
+$\epsilon_{R}=\prod_{i=1}^{k-1}\frac{1-c_{i}^{4}}{1+c_{i}^{4}}.$
 Proof.
@@
 The base case $k=2$ follows directly from Theorem 4.3.
-The condition $v_{2}-v_{1}\neq\operatorname{diam}({\mathbb{S}^{n-1}})$ ensures $r=\left\langle{v_{1},v_{2}}\right\rangle>-1$.
-Assume the result holds for all configurations of size $<k$, and let $c_{1}=r_{1}$. Define
+The condition $v_{2}-v_{1}\neq\operatorname{diam}({\mathbb{S}^{n-1}})$ ensures $r=\left\langle{v_{1},v_{2}}\right\rangle>-1$ and the condition $\sigma(A)\geq\omega_{R}(n^{-\epsilon_{R}})$ ensures that the error term goes to 0.
+Assume the result holds for all configurations of size <k<k, and let $c_{1}=r_{1}$. Define
 $A_{\mathrm{good}}=\left\{x\in{\mathbb{S}^{n-1}}:\sigma_{x,c_{1}}(A\cap S_{x,c_{1}})\geq\frac{1}{2}\sigma(A)^{(1+c_{1}^{4})/(1-c_{1}^{4})}\right\}.$
@@
 Given $x_{1}$, the event $x_{2}\in A,\cdots,x_{k}\in A$ is equal to $y_{2}\in A\cap S_{x_{1},c_{1}},\cdots,y_{k}\in A\cap S_{x_{1},c_{1}}$.
-Conditioned on $x_{1}\in A_{\mathrm{good}}$, we have $\sigma_{x_{1},c_{1}}(A\cap S_{x_{1},c_{1}})\geq\frac{1}{2}\sigma(A)^{(1+c_{1}^{4})/(1-c_{1}^{4})}$, so by the inductive hypothesis:
+Conditioned on $x_{1}\in A_{\mathrm{good}}$, we have $\sigma_{x_{1},c_{1}}(A\cap S_{x_{1},c_{1}})\geq\frac{1}{2}\sigma(A)^{(1+c_{1}^{4})/(1-c_{1}^{4})}\geq\omega_{R^{\prime}}(n^{-\epsilon_{R^{\prime}}})$, so by the inductive hypothesis:
 $\mathop{\bf Pr\/}(y_{2},\dots,y_{k}\in A\cap S_{x_{1},c_{1}})\geq\Omega_{R}\left(\sigma(A)^{C_{R^{\prime}}(1+c_{1}^{4})/(1-c_{1}^{4})}\right).$
@@
 $\mathop{\bf Pr\/}(x_{1}\in A_{\mathrm{good}})\geq\Omega_{R}(\sigma(A)^{2/(1-c_{1}^{4})}).$
-Multiplying the two bounds gives
+Multiplying the two bounds give
 $\mathop{\bf Pr\/}(x_{1},\dots,x_{k}\in A)\geq\Omega_{R}\left(\sigma(A)^{C_{R}}\right),$
@@
 The condition
-$v_{k}-v_{k-1}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right)$
+$\left\lVert v_{k}-v_{k-1}\right\rVert_{2}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right)$
 is equivalent to $c_{k-1}\neq-1$, which is necessary for applying Theorem 4.3.
-For a given configuration $(v_{1},\dots,v_{k})$, the quantity $c_{i}$ corresponds to the inner product between $v_{i}$ and later vectors $v_{j}$, after projecting to the $(n-i)$-subsphere $\bigcap_{l<i}S_{v_{l},r_{l}}$ and normalizing.
-Since $v_{1},\dots,v_{k}$ are distinct, we have $-1<c_{i}<1$ for all $i\leq k-2$ and $c_{k-1}<1$.
+For a given configuration $(v_{1},\dots,v_{k})$, the quantity $c_{i}$ corresponds to the inner product between $v_{i}$ and later vectors $v_{j}$, after projecting to the (ni)(n-i)-subsphere $\bigcap_{l<i}S_{v_{l},r_{l}}$ and normalizing.
+Since $v_{1},\dots,v_{k}$ are distinct, we have $1<c_{i}<1$ for all $i\leq k-2$ and $c_{k-1}<1$.
 The only forbidden case is $c_{k-1}=-1$, which occurs precisely when $v_{k}-v_{k-1}$ equals the diameter of the intersection $\bigcap_{j<k-1}S_{v_{j},r_{j}}$.
@@
 Fix any $r\in\left(-\frac{1}{k-1},1\right)$.
-For any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega(n^{-2/3})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)$ lies entirely in $A$ satisfies
-$\mathop{\bf Pr\/}_{(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)}(x_{1}\in A,\dots,x_{k}\in A)\geq\Omega_{k,r}(\sigma(A)^{C_{R}}),$
-as $n\to\infty$. The constant $C_{R}$ is given by
-$C_{R}=\sum_{i=1}^{k-1}\frac{2}{1-c_{i}^{4}}\prod_{j=1}^{i-1}\frac{1+c_{j}^{4}}{1-c_{j}^{4}},$
+For any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega_{k,r}(n^{-\epsilon_{k,r}})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)$ lies entirely in $A$ satisfies
+$\mathop{\bf Pr\/}_{(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)}(x_{1}\in A,\dots,x_{k}\in A)\geq\Omega_{k,r}(\sigma(A)^{C_{k,r}}),$
+as $n\to\infty$. The constants $C_{k,r},\ \epsilon_{k,r}$ are given by
+$C_{k,r}=\sum_{i=1}^{k-1}\frac{2}{1-c_{i}^{4}}\prod_{j=1}^{i-1}\frac{1+c_{j}^{4}}{1-c_{j}^{4}}\qquad\epsilon_{k,r}=\prod_{i=1}^{k-1}\frac{1-c_{i}^{4}}{1+c_{i}^{4}},$
 where
@@
 5 Open Problems
-Our density version of the Frankl–Rödl theorem on the sphere (Theorem 4.3) requires the set $A\subseteq\mathbb{S}^{n-1}$ to have spherical measure $\sigma(A)>\omega(n^{-2/3})$ in order to guarantee a non-trivial lower bound.
-In contrast, the classical Frankl–Rödl theorem asserts that for any $-1<r<1$, there exists a constant $\epsilon=\epsilon(r)>0$ such that
+Our density version of the Frankl–Rödl theorem on the sphere (Theorem 4.3) requires the set $A\subseteq\mathbb{S}^{n-1}$ to have spherical measure $\sigma(A)>\omega\left(\frac{1}{n}\right)$ in order to guarantee a non-trivial lower bound.
+In contrast, the classical Frankl–Rödl theorem asserts that for any 1<r<1-1<r<1, there exists a constant $\epsilon=\epsilon(r)>0$ such that
 $\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,y\in A)>0$
@@
 $\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,y\in A)$
-for all $-1<r<1$, even when $\sigma(A)$ is exponentially small in $n$.
+for all 1<r<1-1<r<1, even when $\sigma(A)$ is exponentially small in $n$.
 Another direction concerns the class of configurations we consider.
-Our density result applies to inductive configurations, but Matoušek and Rödl [MR95] showed that for any configuration $P$ with circumradius less than 1, the probability
-$\mathop{\bf Pr\/}_{(x_{1},\ldots,x_{k})\in P}(x_{i}\in A\ \forall i)>0$
-is positive for any set $A\subseteq\mathbb{S}^{n-1}$ with constant density $\sigma(A)>\epsilon(P)$.
-It remains open whether one can establish a density version of this theorem—that is, to derive quantitative lower bounds on the probability in terms of $\sigma(A)$.
+Our density result applies to inductive configurations, but Matoušek and Rödl [MR95] showed that fix constant $c$ and for any configuration $P$ with circumradius less than 1, any $c$ coloring of $\mathbb{S}^{n-1}$} contains a monochromatic copy of $P$.
+It remains open whether one can establish a density version of this theorem—that is, to show
+$\mathop{\bf Pr\/}_{(x_{1},\ldots,x_{k})\in P}(x_{i}\in A\ \forall i)>\epsilon(\sigma(A))$
+for any set $A\subseteq\mathbb{S}^{n-1}$ with constant density $\sigma(A)>\epsilon(P)$ where $\epsilon(\sigma(A))$ is a lower bound on the probability in terms of $\sigma(A)$.
 Finally, our current results can be almost extended to 3-point configurations. For example, using Theorem 4.3, we can obtain bounds on